# TorchTitan: One-stop PyTorch native solution for production ready LLM pre-training

### v2

@@ Abstract
This paper introduces TorchTitan, an open-source, PyTorch-native distributed training system that unifies and advances state-of-the-art techniques, streamlining integration and reducing engineering overhead. TorchTitan enables seamless application of 3D parallelism in a modular and composable manner, while featuring elastic scaling to adapt to changing computational requirements. The system provides comprehensive logging, efficient checkpointing, and debugging tools, ensuring production-ready training. Moreover, TorchTitan incorporates innovative hardware-software co-designed solutions, leveraging cutting-edge features like Float8 training and SymmetricMemory to maximize hardware utilization. As a flexible experimental test bed, TorchTitan facilitates the curation and comparison of custom recipes for diverse training contexts. By leveraging TorchTitan, we developed optimized training recipes for the Llama 3.1 family and provide actionable guidance on selecting and combining distributed training techniques to maximize training efficiency, based on our hands-on experiences.
We thoroughly assess TorchTitan on the Llama 3.1 family of LLMs, spanning 8 billion to 405 billion parameters, and showcase its exceptional performance, modular composability, and elastic scalability. By stacking training optimizations, we demonstrate accelerations of 65.08% with 1D parallelism at the 128-GPU scale (Llama 3.1 8B), an additional 12.59% with 2D parallelism at the 256-GPU scale (Llama 3.1 70B), and an additional 30% with 3D parallelism at the 512-GPU scale (Llama 3.1 405B) on NVIDIA H100 GPUs over optimized baselines.
@@ 1 Introduction
LLMs are at the forefront of NLP advancement. Large Language Models (LLMs) (Devlin, 2018; Liu et al., 2019; Radford et al., 2019; Chowdhery et al., 2023; Anil et al., 2023; Achiam et al., 2023; Dubey et al., 2024; Jiang et al., 2024; Abdin et al., 2024) have been driving force behind the advancement of natural language processing (NLP) applications spanning language translation, content/code generation, conversational AI, text data analysis, creative writing and art, education and research etc.
LLMs require billions of parameters and training over trillion tokens to achieve state-of-the-art performance. Achieving state-of-the-art LLM performance requires massive scale, exemplified by top-performing models like Llama 3.1 (405B parameters, 15T tokens, 30.84M GPU hours, 16K H100 GPUs) (Dubey et al., 2024) and Google’s PaLM (540B parameters, 0.8T tokens, 9.4M TPU hours, 6144 TPUv4 chips) (Chowdhery et al., 2023). These models demonstrate exceptional natural language understanding and generation capabilities, but necessitate substantial computational resources, memory, and time to train, highlighting the significant investment required to advance natural language processing.
LLM training challenges are being tackled from all sides. Training large language models (LLMs) at scale is a daunting task that requires a delicate balance of parallelism, computation, and communication, all while navigating intricate memory and computation tradeoffs. The massive resources required for training make it prone to GPU failures, underscoring the need for efficient recovery mechanisms and checkpointing strategies to minimize downtime (Eisenman et al., 2022; Wang et al., 2023; Gupta et al., 2024; Maurya et al., 2024; Wan et al., 2024). To optimize resource utilization and achieve elastic scalability, it is crucial to combine multiple parallelism techniques, including Data Parallel (Li et al., 2020; Rajbhandari et al., 2020; Zhang et al., 2022; Zhao et al., 2023), Tensor Parallel (Narayanan et al., 2021; Wang et al., 2022; Korthikanti et al., 2023), Context Parallel (Liu et al., 2023; Liu and Abbeel, 2024; NVIDIA, 2023; Fang and Zhao, 2024), and Pipeline Parallel (Huang et al., 2019; Narayanan et al., 2019, 2021; Qi et al., 2023). By stacking these parallelisms with memory and computation optimization techniques, such as activation recomputation (Chen et al., 2016; Korthikanti et al., 2023; He and Yu, 2023), mixed precision training (Micikevicius et al., 2018, 2022), and deep learning compilers (Bradbury et al., 2018; Yu et al., 2023; Li et al., 2024; Ansel et al., 2024), it is possible to maximize hardware utilization.
Limitations of existing systems incorporating state-of-the-art techniques. While state-of-the-art distributed training techniques have significantly advanced the field, existing systems that incorporate them still fall short in addressing critical challenges that hinder their usability, adoption and effectiveness for researchers and industry practitioners.
@@
    Non-composable: Existing systems struggle to combine and stack various parallelism techniques, limiting the exploration of multi-dimensional parallelism. Further integrating them with memory and computation optimizations is challenging, hindering training efficiency.
@@
    Inflexible and monolithic architecture: Current systems are not modular or extensible, making it difficult to integrate and compare new techniques, optimizations, and hardware, and limiting adaptability to evolving machine learning landscapes.
@@
    Inefficient hardware utilization: Current systems fail to fully leverage advanced hardware features, leading to sub-optimal GPU efficiency, and lack customizable activation checkpointing strategies to navigate memory-computation trade-offs.
@@
    Insufficient support for production-grade training: Existing systems lack scalable and efficient distributed checkpointing, making failure recovery and model saving cumbersome, and often do not provide adequate debugging tools and logging metrics, leading to difficulties in identifying and fixing issues, particularly for those without extensive expertise.
@@
    Existing systems fall short in harnessing the full potential of frameworks like PyTorch, missing out on bug fixes, optimized kernels, new features, and compiler support. They also rely on external dependencies that often lack thorough testing and can become outdated or incompatible due to inadequate maintenance.
Root cause: Lack of an expressive tensor abstraction. The root cause of non-composability and inflexibility of a distributed system stems from the lack of using an expressive tensor and device abstraction as a central component, upon which all of the distributed parallelisms, checkpointing, and efficiency optimizations can be built.
Design Principle: Unified distributed tensor and device abstractions as building blocks. A unified device abstraction represents the distributed system as a multi-dimensional array, where each dimension corresponds to a parallelism technique, managing communication between devices and handling collective process groups. A complementary tensor abstraction enables tensors to be sharded across this array, maintaining sharding specifications and supporting automatic sharding propagation. Together, these abstractions enable seamless composition of parallelism techniques, ensure correct semantics, and facilitate dispatching of collectives for distributed operations
We address the technical challenge of a unified tensor abstraction by employing PyTorch’s Distributed Tensor (DTensor) and DeviceMesh (Wanchao Liang, 2023) as the foundational components for TorchTitan. Through our work with DTensor and DeviceMesh, we identified key limitations and addressed them. By using and extending DTensor, we develop TorchTitan, a production-ready system that enables composability, modularity, flexibility, and extensibility in distributed training. TorchTitan facilitates the composition of 3D parallelism, training optimizations, scalable distributed checkpointing, and harnesses the full benefits of the PyTorch ecosystem.
@@
    We demonstrate how to compose and stack various parallelism techniques, facilitating the exploration of multi-dimensional parallelism in large language model training (§2.1).
@@
    We offer production grade training by incorporating scalable and efficient distributed checkpointing to facilitate fast failure recovery, integrating debugging tools like Flight Recorder to debug crashed/stuck jobs, and providing extensive logging metrics (§2.3).
@@
    We extensively evaluate TorchTitan on the Llama 3.1 family of models (8B, 70B, and 405B with 1D, 2D, and 3D parallelisms, respectively) at the scale from 8 to 512 GPUs to demonstrate elastic scalability while ensuring efficiency, convergence, and accuracy. In summary, we demonstrate training accelerations of 65.08% with 1D parallelism at the 128-GPU scale (Llama 3.1 8B), an additional 12.59% with 2D parallelism at the 256-GPU scale (Llama 3.1 70B), and an additional 30% with 3D parallelism at the 512-GPU scale (Llama 3.1 405B) on latest NVIDIA H100 GPUs over optimized baselines (§3.2).
@@
  7. 7.
    We show how our modular and extensible architecture allows for seamless integration and comparison of new techniques, optimizations, and hardware, ensuring adaptability to evolving machine learning landscapes (§4).
By providing an accessible and extensible platform, TorchTitan democratizes large language model (LLM) pre-training, empowering a wider range of researchers and developers to tap into the potential of LLMs and accelerate innovation in the field.
@@ 2 Elasticity through Composability
TorchTitan incorporates various parallelism techniques in a modular manner to enable easy, user-selectable combinations of multi-dimensional parallelisms. This composability enables the tackling of difficult scaling challenges by enhancing the ease of frontier exploration for optimizing training efficiencies at scale.
The codebase of TorchTitan is organized purposefully to enable composability and extensibility. We intentionally keep three main components separate and as orthogonal as possible: (1) the model definition, which is parallelism-agnostic and designed for readability, (2) parallelism helpers, which apply Data Parallel, Tensor Parallel, and Pipeline Parallel to a particular model, and (3) a generalized training loop. All these components are configurable via TOML files with command-line overrides, and it is easy to add new models and parallelism techniques on top of the existing codebase.
@@ 2.1 Composable N-D parallelism training
In this section, we will walk through the entire regime of scaling model training on large clusters, including meta device initialization and the core composable multi-dimensional parallelisms, to showcase how these techniques can be composed to train LLMs efficiently at increasing scale in TorchTitan. The corresponding actual code snippets in TorchTitan can be found in Appendix 8.
@@ 2.1.1 Large-scale model initialization using meta device
Given the exponential increase in model sizes for LLMs, the first scaling issue appears even before the actual training starts. This is the need to instantiate a large model for sharding across the cluster, yet without overflowing CPU or GPU memory.
To tackle this, we enabled meta device initialization for models in TorchTitan, where the model is first initialized on a “meta” device type. The meta device tensor only holds the metadata information, not the actual data, making initialization ultra-fast. After that, we perform model sharding and transforming the model parameters into Distributed Tensors (DTensors) where each parameter holds a local shard that lives on the meta device. Finally, we perform parameter initialization based on the user-defined initialization functions. We leverage Distributed Tensor to properly sync Random Number Generator (RNG) seeds, and initialize the parameters according to their sharding layouts. This ensures the parameters start with the same values as if the whole model were initialized on one device before sharding, and thus facilitating convergence comparisons between different parallelism configurations.
@@ 2.1.2 Fully Sharded Data Parallel
The original Fully Sharded Data Parallel (FSDP) (Zhao et al., 2023) is an effective implementation of ZeRO that offers large model training capability in PyTorch. However, the original implementation (FSDP1) in PyTorch suffers from various limitations due to its FlatParameter implementation (see details in Appendix 9.1).
Given these limitations, TorchTitan integrates a new version of Fully Sharded Data Parallel (FSDP2), which uses the per-parameter Distributed Tensor sharding representation and thus provides better composability with model parallelism techniques and other features that require the manipulation of individual parameters,
@@
For scaling to even larger world sizes, TorchTitan also integrates Hybrid Sharded Data Parallel (HSDP) which extends FSDP2 by creating sharding groups. Details are shown in Appendix 9.2
@@ 2.1.3 Tensor Parallel
TP is implemented in TorchTitan using the PyTorch’s RowwiseParallel and ColwiseParallel APIs, where the model parameters are partitioned to DTensors and perform sharded computation with it. By leveraging DTensor, the TP implementation does not need to touch the model code, which allows faster enablement on different models and provides better composability with other features mentioned in this paper.
@@
While TP partitions the most computationally demanding aspects, Sequence Parallel (SP) perform sharded computation for the normalization or dropout layers on the sequence dimension, which otherwise generate large replicated activation tensors, and thus can be challenging to memory constraints per GPU. See Appendix 9.3 for more details, illustrations, and usage for both TP and FSDP + TP.
Because of the synergistic relationship between TP and SP, TorchTitan natively bundles these two together and they are jointly controlled by the TP degree setting.
@@
When the loss function is computed, the model outputs are usually very large. Since the model outputs from TP/SP are sharded on the (often huge) vocabulary dimension, naively computing the cross-entropy loss requires gathering all the shards along the TP dimension to make the outputs be replicated, which incurs large memory usage.
With Loss Parallel, the cross entropy loss can be computed efficiently, without gathering all the model output shards to every single GPU. This not only significantly reduces the memory consumption, but also improves training speed by reducing communication overhead and doing sharded computation in parallel. Given these improvements, TorchTitan implements loss parallel by default.
@@ 2.1.4 Pipeline Parallel
For pre-training at the largest scales, TorchTitan offers Pipeline Parallelism, which becomes essential due to having the lightest amount of communication overhead and leveraging P2P communications.
Pipeline Parallel (PP) views the model as a sequence of operations, chunking the operations (and the parameters used by them) into S stages which run on separate groups of devices. In the typical case, one stage represents a single model layer or a group of N adjacent model layers, but in theory it could be even be a partial layer. For the forward pass, a stage receives input activations (except stage 0), performs local computation, and sends output activations (except stage S - 1). The last stage performs a loss computation, and begins the backward pass, sending gradients in the reverse order through the pipeline. To improve efficiency, the input batch is broken into microbatches and a pipeline schedule overlaps computation on one microbatch with communication for others. TorchTitan enables a variety of pipeline schedules, with their schedules previously described in other works (Narayanan et al., 2019; Huang et al., 2019; Narayanan et al., 2021; Qi et al., 2023).
The training loop must also account for creation of pipeline stages, and executing a pipeline schedule rather than invoking model.forward() directly. Because the schedule computes loss per microbatch, the loss computation and any logging code must be updated for PP. In TorchTitan, we propose to define a shared loss_fn to be used by both pipeline and non-pipeline code paths, and thus minimise divergence in the training loop.
Interactions with data parallelism, such as ensuring that data-parallel reduction happens only after the last microbatch in the schedule, and scheduling of shard and unshard operations when using zero-3, are also handled transparently inside the pipeline schedule executor, simplifying the trainer implementation in TorchTitan. For it’s usage in TorchTitan, please see Appendix 9.4.

### v3

# TorchTitan: One-stop PyTorch native solution for production ready LLM pretraining

@@ Abstract
This paper introduces TorchTitan, an open-source, PyTorch-native distributed training system that unifies and advances state-of-the-art techniques, streamlining integration and reducing engineering overhead. TorchTitan enables seamless application of 4D parallelism in a modular and composable manner, while featuring elastic scaling to adapt to changing computational requirements. The system provides comprehensive logging, efficient checkpointing, and debugging tools, ensuring production-ready training. Moreover, TorchTitan incorporates innovative hardware-software co-designed solutions, leveraging cutting-edge features like Float8 training and SymmetricMemory to maximize hardware utilization. As a flexible experimental test bed, TorchTitan facilitates the curation and comparison of custom recipes for diverse training contexts. By leveraging TorchTitan, we developed optimized training recipes for the Llama 3.1 family and provide actionable guidance on selecting and combining distributed training techniques to maximize training efficiency, based on our hands-on experiences.
We thoroughly assess TorchTitan on the Llama 3.1 family of LLMs, spanning 8 billion to 405 billion parameters, and showcase its exceptional performance, modular composability, and elastic scalability. By stacking training optimizations, we demonstrate accelerations ranging from 65.08% on Llama 3.1 8B at 128 GPU scale (1D), 12.59% on Llama 3.1 70B at 256 GPU scale (2D), to 30% on Llama 3.1 405B at 512 GPU scale (3D) on NVIDIA H100 GPUs over optimized baselines. We also demonstrate the effectiveness of 4D parallelism in enabling long context training.
@@ 1 Introduction
Large Language Models (LLMs) (Devlin, 2018; Liu et al., 2019; Radford et al., 2019; Chowdhery et al., 2023; Anil et al., 2023; Achiam et al., 2023; Dubey et al., 2024; Jiang et al., 2024; Abdin et al., 2024) have been the driving force behind the advancement of natural language processing (NLP) applications spanning language translation, content/code generation, conversational AI, text data analysis, creative writing and art, education, and research, etc.
Achieving state-of-the-art LLM performance requires massive scale, exemplified by top-performing models like Llama 3.1 (405B parameters, 15T tokens, 30.84M GPU hours, 16K H100 GPUs) (Dubey et al., 2024) and Google’s PaLM (540B parameters, 0.8T tokens, 9.4M TPU hours, 6144 TPUv4 chips) (Chowdhery et al., 2023). These models demonstrate exceptional natural language understanding and generation capabilities, but at the same time necessitate substantial computational resources, memory, and time to train, highlighting the significant investment required to advance natural language processing.
Training large language models (LLMs) at scale is a daunting task that requires a delicate balance of parallelism, computation, and communication, all while navigating intricate memory and computation trade-offs. The massive resources required for training make it prone to GPU failures, underscoring the need for efficient recovery mechanisms and checkpointing strategies to minimize downtime (Eisenman et al., 2022; Wang et al., 2023; Gupta et al., 2024; Maurya et al., 2024; Wan et al., 2024). To optimize resource utilization and achieve elastic scalability, it is crucial to combine multiple parallelism techniques, including Data Parallel (Li et al., 2020; Rajbhandari et al., 2020; Zhang et al., 2022; Zhao et al., 2023), Tensor Parallel (Narayanan et al., 2021; Wang et al., 2022; Korthikanti et al., 2023), Context Parallel (Liu et al., 2023; Liu and Abbeel, 2024; NVIDIA, 2023; Fang and Zhao, 2024), and Pipeline Parallel (Huang et al., 2019; Narayanan et al., 2019, 2021; Qi et al., 2023). By stacking these parallelisms with memory and computation optimization techniques, such as activation recomputation (Chen et al., 2016; Korthikanti et al., 2023; He and Yu, 2023; Purandare et al., 2023), mixed precision training (Micikevicius et al., 2018, 2022), and deep learning compilers (Bradbury et al., 2018; Yu et al., 2023; Li et al., 2024; Ansel et al., 2024), it is possible to maximize hardware utilization.
While state-of-the-art distributed training techniques have significantly advanced the field, existing systems that incorporate them still fall short in addressing critical challenges that hinder their usability, adoption and effectiveness for researchers and industry practitioners.
@@
    Non-composable: Existing systems struggle to integrate and stack parallelism techniques, limiting multi-dimensional exploration and integration with memory and computation optimizations, thereby reducing training efficiency.
@@
    Inflexible Architecture: Lack of modularity and extensibility hampers the integration of new techniques, optimizations, and hardware, limiting adaptability to evolving ML landscapes.
@@
    Inefficient Hardware Utilization: Poor leverage of advanced hardware features results in sub-optimal GPU efficiency and lack of customizable checkpointing strategies for memory-computation trade-offs.
@@
    Insufficient Support for Production Training: Limited distributed checkpointing scalability, cumbersome failure recovery, and inadequate debugging tools hinder production-grade workflows.
@@
    Framework Limitations: Dependence on external, poorly maintained dependencies and failure to harness PyTorch’s optimized kernels, new features, and compiler support lead to inefficiencies and compatibility issues.
The non-composability and inflexibility of distributed systems stem from the absence of unified tensor and device abstractions applied consistently across the stack. Without these foundational components, parallelism strategies, checkpointing, and efficiency optimizations remain fragmented, limiting modularity, scalability, and extensibility.
TorchTitan ’s primary research contribution lies in identifying and unifying the core principles of parallelism and optimization techniques into a cohesive framework. By leveraging and extending PyTorch’s Distributed Tensor (DTensor) and DeviceMesh (Wanchao Liang, 2023), TorchTitan provides a unified abstraction that simplifies the composition of parallelism strategies, and ensures correct single device semantics with its sharding primitives. Unlike existing systems that often rely on rigid or ad-hoc designs, TorchTitan introduces a unified template for distributed training, enabling researchers to systematically explore configurations, rigorously evaluate existing methods, and uncover novel techniques within the design space.
TorchTitan represents a complete distributed training system for large language models (LLMs), rather than merely a collection of individual techniques. Its modular, extensible architecture supports seamless composition of 4D parallelism, advanced training optimizations, and scalable distributed checkpoint save/load, all while harnessing PyTorch’s native capabilities. The system not only enable production-grade training with thousands of GPUs, but also reduces complexity and fosters innovation, setting a new standard for scalable and flexible distributed training systems.
@@
    We demonstrate how to compose various parallelism techniques, facilitating the exploration of multi-dimensional parallelism in large language model training (§2.1).
@@
    We offer production grade training by incorporating scalable and efficient distributed checkpoint to facilitate fast failure recovery, integrating debugging tools like Flight Recorder to debug crashed/stuck jobs, and provide extensive logging metrics (§2.3).
@@
    We extensively evaluate TorchTitan on Llama 3.1 family of models, stacking 1D to 4D parallelisms (respectively), at the scale from 8 to 512 GPUs to demonstrate elastic scalability while ensuring efficiency, convergence, and accuracy. In summary, we demonstrate training accelerations ranging from 65.08% on Llama 3.1 8B at 128 GPU scale (1D), 12.59% on Llama3.1 70B at 256 GPU scale (2D), to 30% on Llama3.1 405B at 512 GPU scale (3D), and the effectiveness of 4D parallelism in enabling long context training, on latest NVIDIA H100 GPUs over optimized baselines (§3.2).
@@
By providing an accessible and extensible platform, TorchTitan democratizes large language model (LLM) pretraining, empowering a wider range of researchers and developers to tap into the potential of LLMs and accelerate innovation in the field.
@@ 2 Elasticity through composability
TorchTitan incorporates various parallelisms in a modular manner to enable easy, user-selectable combinations of multi-dimensional shardings. This composability enables the tackling of difficult scaling challenges by enhancing the ease of exploration for optimizing training efficiencies at scale.
The codebase of TorchTitan is organized purposefully to enable composability and extensibility. We intentionally keep three main components separate and as orthogonal as possible: (1) the model definition, which is parallelism-agnostic and designed for readability, (2) parallelism helpers, which apply parallelisms and training optimizations to a particular model, and (3) a generalized training loop. All these components are configurable via TOML files with command-line overrides, and it is easy to add new models and parallelism techniques on top of the existing codebase.
@@ 2.1 Composable N-D parallelism training
In this section, we will walk through the entire regime of scaling model training on large clusters, including meta device initialization and the core composable multi-dimensional parallelisms, to showcase how these techniques can be composed to train LLMs efficiently at increasing scale in TorchTitan. The corresponding code snippets in TorchTitan can be found in Appendix 7.
@@ 2.1.1 Large-scale model initialization using meta device
As LLMs grow exponentially, scaling challenges arise even before training begins, particularly in instantiating large models for sharding without exceeding CPU or GPU memory limits.
To address this, TorchTitan enables meta device initialization, where the model is first created on a meta device that stores only metadata, making initialization ultra-fast. The model is then sharded into Distributed Tensors (DTensors), with the local shard of each parameter residing on the meta device. Finally, parameter initialization is performed using user-defined functions, ensuring correct DTensor sharding layouts and proper RNG seed usage.
@@ 2.1.2 Fully Sharded Data Parallel
The original Fully Sharded Data Parallel (FSDP) (Zhao et al., 2023) is an effective implementation of ZeRO that offers large model training capability in PyTorch. However, the original implementation (FSDP1) in PyTorch suffers from various limitations due to its FlatParameter implementation.
Given these limitations, TorchTitan integrates a new version of Fully Sharded Data Parallel (FSDP2), which uses the per-parameter Distributed Tensor sharding representation and thus provides better composability with model parallelism techniques and other features that require the manipulation of individual parameters.
@@
For scaling to even larger world sizes, TorchTitan also integrates Hybrid Sharded Data Parallel (HSDP) which extends FSDP2 by creating 2D DeviceMesh with replica groups. Details are shown in Appendix 8.2
@@ 2.1.3 Tensor Parallel
TP is implemented in TorchTitan using the PyTorch’s RowwiseParallel and ColwiseParallel APIs, where the model parameters are partitioned to DTensors and perform sharded computation with it (Figure 3). By leveraging DTensor, the TP implementation does not need to touch the model code, which allows faster enablement on different models and provides better composability with other features mentioned in this paper.
@@
While TP partitions the most computationally demanding aspects, Sequence Parallel (SP) performs a sharded computation for the normalization or dropout layers on the sequence dimension, which otherwise generate large replicated activation tensors, and thus can be challenging to memory constraints per GPU. See Appendix 8.3 for more details, illustrations, and usage for both TP and FSDP + TP.
Due to the synergistic relationship between TP and SP, TorchTitan natively bundles these two together, and they are jointly controlled by the TP degree setting.
@@
When computing the loss function, model outputs are typically large, especially with TP/SP, where they are sharded across the vocabulary dimension. Naively computing cross-entropy loss requires gathering all shards, leading to high memory usage.
Loss Parallel enables efficient loss computation without fully gathering model outputs, significantly reducing memory consumption and improving training speed by minimizing communication overhead and enabling parallel sharded computation. Due to these advantages, TorchTitan implements Loss Parallel by default.
@@ 2.1.4 Pipeline Parallel
For large-scale pretraining, TorchTitan employs Pipeline Parallelism (PP), which minimizes communication overhead by leveraging P2P communications. PP divides the model into $S$ stages, each running on a separate group of devices. Typically, each stage represents a model layer or a group of adjacent layers, but can include partial layers. During the forward pass, each stage receives input activations (except stage 0), computes locally, and sends output activations (except stage $S-1$). The last stage computes the loss and initiates the backward pass, sending gradients in reverse order. To improve efficiency, the input batch is split into microbatches, and the pipeline schedule overlaps computation and communication across microbatches. TorchTitan supports various pipeline schedules (Narayanan et al., 2019; Huang et al., 2019; Narayanan et al., 2021; Qi et al., 2023). Recently, TorchTitan added support for new schedules including ZeroBubble and ’Flexible-Interleaved-1F1B’, making use of pipeline IR to quickly express new schedules as a list of compute actions and rely on compiler passes to insert and optimize communication actions PyTorch Team 2024d.
The PP training loop differs from standard training by creating pipeline stages and executing schedules instead of directly invoking model.forward(). Since loss is computed per microbatch, TorchTitan introduces a shared loss_fn to unify pipeline and non-pipeline workflows, reducing code divergence.
torch.distributed.pipelining also simplifies interactions with data parallelism, ensuring that reductions occur only after the final microbatch and handling shard/unshard operations (e.g., with ZeRO-3), as well as applying gradient scaling transparently within the pipeline schedule executor. For more details on TorchTitan’s implementation of PP, see Appendix 8.4.
2.1.5 Context Parallelism
TorchTitan has been extended to incorporate Context Parallelism (CP) (Liu et al., 2023; Liu and Abbeel, 2024; NVIDIA, 2023), enabling 4D parallelism by adding CP as an additional dimension to existing DP, TP, and PP. CP scales model training by splitting the sequence dimension across GPUs, significantly increasing the maximum trainable context length without causing out-of-memory (OOM) errors. For example, on Llama 3.1 8B with 8 H100 GPUs, using CP enabled training at context lengths up to 262,144 tokens, achieving minor MFU degradation as CP degree increases (PyTorch Team, 2025). For more details on CP integration please refer to Appendix 8.5.